OpenStudy (amonoconnor):

How do you simply the following expression: (((sqrt(6-x))-2)/((sqrt(3-x)-1)))*(((sqrt(3-x))+1)/((sqrt(3-x))+1))) Thank you very much! Any and all help is greatly appreciated!

OpenStudy (mertsj):

\[\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1}\times \frac{\sqrt{3-x}+1}{\sqrt{3-x}+1}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (amonoconnor):

Yes!

OpenStudy (mertsj):

In the denominator of the first fraction: is it -1or +1

OpenStudy (jackthegreatest):

well for one u can cross out the equation on the right, the top and bottom r exactly the same

OpenStudy (amonoconnor):

Shoot.. you are correct. "-1"

OpenStudy (mertsj):

If the problem is as I have posted it the denominators are of the form (a-b)(a+b) and that is equal to a^2-b^2

OpenStudy (amonoconnor):

I know, that's the point: Multiplying by the conjugate, but I'm weary as to how to multiple that across... ?

OpenStudy (mertsj):

So the denominator is 3-x-1or 2-x

OpenStudy (mertsj):

now let's do the numerator multiplication:

OpenStudy (amonoconnor):

How did you get that? :/ I'm not disagreeing, I just don't know how you worked it out..

OpenStudy (mertsj):

\[(\sqrt{6-x}-2)(\sqrt{3-x}+1)=\sqrt{(6-x)(3-x)}+\sqrt{6-x}-2\sqrt{3-x}-2\]

OpenStudy (mertsj):

What is (a-b)(a+b)

OpenStudy (amonoconnor):

(18 - 9x + x) :)

OpenStudy (mertsj):

Where did that come from?

OpenStudy (amonoconnor):

Isn't that what'll be under the radical, in the first term in the numerator? :/

OpenStudy (mertsj):

It would be 18-9x+x^2

OpenStudy (amonoconnor):

Dude, I'm tired.. my bad again.

OpenStudy (mertsj):

so put that numerator over 2-x and you are done.

OpenStudy (amonoconnor):

Can we cancel out anything else?

OpenStudy (mertsj):

Are there any common factors?

OpenStudy (amonoconnor):